OpenStudy (anonymous):

Show how to solve 8x^2+14x+5=0 by completing the square.

OpenStudy (alexandervonhumboldt2):

8x^2+14x+5=0 Multiply 8 by 5. we get 40. Which two numbers add up to 14 and multiply to 40? They are 10 and 4. rewrite 14x as the sum of 10x and 4x: 8x2+10x+4x+5=0 2x(4x+5)+(4x+5)=0 (4x+5)(2x+1)=0 now find the roots

OpenStudy (alexandervonhumboldt2):

use the viet theorem

OpenStudy (anonymous):

thank you @AlexandervonHumboldt2